# Boundary-driven quantum systems near the Zeno limit: steady states and long-time behavior

### v1

@@ Abstract
We study composite open quantum systems with a finite-dimensional state space $\mathcal{H}_{AB}=\mathcal{H}_{A}\otimes\mathcal{H}_{B}$ governed by a Lindblad equation ${\frac{{\rm d}}{{\rm d}t}\rho(t)=\mathcal{L}_{\gamma}\rho(t)}$ where $\mathcal{L}_{\gamma}\rho=-i[H,\rho]+\gamma\mathcal{D}\rho$. Here, $H$ is a Hamiltonian on $\mathcal{H}_{AB}$ while $\mathcal{D}$ is a dissipator $\mathcal{D}_{A}\otimes\mathds{1}$ acting non-trivially only on part $A$ of the system, which can be thought of as the boundary, and $\gamma$ is a parameter. It is known that the dynamics simplifies as the Zeno limit, $\gamma\to\infty$, is approached: after a initial time of order $\gamma^{-1}$, $\rho(t)$ is well approximated by $\pi_{A}\otimes R(t)$ where $\pi_{A}$ is a density matrix on $\mathcal{H}_{A}$ such that $\mathcal{D}_{A}\pi_{A}=0$, and $R(t)$ is an approximate solution of $\frac{{\rm d}}{{\rm d}t}R(t)=\mathcal{L}_{P,\gamma}R(t)$} where $\mathcal{L}_{P,\gamma}R:=-i[H_{P},R]+\gamma^{-1}\mathcal{D}_{P}R$ with $H_{P}$ being a Hamiltonian on $\mathcal{H}_{B}$ and $\mathcal{D}_{P}$ being a Lindblad generator acting on density matrices on $\mathcal{H}_{B}$. We give a rigorous proof of this holding in greater generality than in previous work; we assume only that $\mathcal{D}_{A}$ is ergodic and gapped. Moreover, we precisely control the error terms, and use this to show that the mixing times of $\mathcal{L}_{\gamma}$ and $\mathcal{L}_{P,\gamma}$ are tightly related near the Zeno limit. Despite this connection, the errors in the approximate description of the evolution accumulate on times of order $\gamma^{2}$, so it is difficult to directly access steady states $\bar{\rho}_{\gamma}$ of $\mathcal{L}_{\gamma}$ through study of $\mathcal{L}_{P,\gamma}$. In order to better control the long time behavior, and in particular the steady states $\bar{\rho}_{\gamma}$, we introduce a third Lindblad generator $\mathcal{D}_{P}^{\sharp}$ that does not involve $\gamma$, but is still closely related to $\mathcal{L}_{\gamma}$ and $\mathcal{L}_{P,\gamma}$. We show that if $\mathcal{D}_{P}^{\sharp}$ is ergodic and gapped, then so are $\mathcal{L}_{\gamma}$ and $\mathcal{L}_{P,\gamma}$ for all large $\gamma$. In this case, if $\bar{\rho}_{\gamma}$ denotes the unique steady state for $\mathcal{L}_{\gamma}$, then $\lim_{\gamma\to\infty}\bar{\rho}_{\gamma}=\pi_{A}\otimes\bar{R}$ where $\bar{R}$ is the unique steady state for $\mathcal{D}_{P}^{\sharp}$. We further show that there is a trace norm convergent expansion $\bar{\rho}_{\gamma}=\pi_{A}\otimes\bar{R}+\gamma^{-1}\sum_{k=0}^{\infty}\gamma^{-k}\bar{n}_{k}$} where, defining $\bar{n}_{-1}:=\pi_{A}\otimes\bar{R}$, $\mathcal{D}\bar{n}_{k}=-i[H,\bar{n}_{k-1}]$ for all $k\geqslant 0$. Using properties of $\mathcal{D}_{P}$ and $\mathcal{D}_{P}^{\sharp}$, we show that this system of equations has a unique solution, and prove convergence. This is illustrated in a simple example for which one can solve for $\bar{\rho}_{\gamma}$, and can carry out the expansion explicitly.
@@ 1 Introduction
$\mathcal{D}:=\mathcal{D}_{A}\otimes\mathds{1}_{B}~,$ (3)
@@
Without loss of generality, a constant has been added to the Hamiltonian so that $\textnormal{Tr}H=0$ , $H=H_{A}\otimes\mathds{1}_{B}+H_{AB}+\mathds{1}_{A}\otimes H_{B}$ ; here $\mathds{1}_{A}$ and $\mathds{1}_{B}$ are the identity operators on those subspaces, $H_{A}=(1/d_{B})\textnormal{Tr}_{B}H$, $H_{B}=(1/d_{A})\textnormal{Tr}_{A}H$, and $d_{A}$ and $d_{B}$ are the dimensions of those subspaces.
@@
uniformly in $t\geqslant t_{\gamma}$ ,
where $R(t)=\textnormal{Tr}_{A}\rho(t)$ is a density matrix on $\mathcal{H}_{B}$. That is, solutions of (1) for a general initial state $\rho_{0}$ and large $\gamma$ rapidly approach $\mathcal{M}$ and then stay close to $\mathcal{M}$, uniformly in time. In Section 3 we give simple proofs of results of this genre in which “closeness” is measured in terms of the trace norm, denoted by $\|\cdot\|_{1}$; see Section 2. In particular, we prove in Theorem 3.1 that for all $\rho_{0}=\pi_{A}\otimes R_{0}\in\mathcal{M}$, there is a constant $C>0$ independent of $\gamma$ such that for all $t>0$,
@@
Note that in (8), the projector $\mathcal{P}$ follows $e^{t\mathcal{L}_{\gamma}}$ while in (7), $\mathcal{P}$ preceded $e^{t\mathcal{L}_{\gamma}}$. These complementary results from Section 3 are important for the study of approximate equations of motion, to which we now turn.
@@
$\lim_{\gamma\to\infty}\|e^{t\mathcal{L}_{\gamma}}\rho_{0}-\pi_{A}\otimes e^{t\mathcal{L}_{P,\gamma}}R_{0}\|_{1}=0\ ,$ (18)
but since for large $\gamma$, $\mathcal{L}_{P,\gamma}$ does not differ much from $\mathcal{K}_{P}$, it follows that (see Theorem 5.3)
@@
In Section 7 we study $e^{-t\mathcal{K}_{P}}e^{t\mathcal{L}_{P,\gamma}}R_{0}$ in the Zeno limit, $\gamma\to\infty$ for general initial data $R_{0}$. We run the initial data forward under the dissipative time evolution governed by $\mathcal{L}_{P,\gamma}$ , and then backwards under the coherent evolution generated by $\mathcal{K}_{P}$ . In the language of scattering theory, we are passing to the “interaction picture” to compare the coherent evolution and its perturbation by a weak dissipation. This is what Davies [13] did in his derivation of Lindblad equations in the weak coupling limit, and hence it is no surprise that his averaging operation introduced there shows up again here.
@@
This has the consequence that if $\mathcal{D}_{P}^{\sharp}$ is ergodic and gapped, then so are $\mathcal{L}_{\gamma}$ and $\mathcal{L}_{P,\gamma}$ for all sufficiently large $\gamma$. See Theorem 7.5 for a precise statement.
This shows one way in which $\mathcal{D}_{P}^{\sharp}$ governs the long-time behavior of the evolution described by both $\mathcal{L}_{P,\gamma}$ , and d by $\mathcal{L}_{\gamma}$ , near the Zeno limit. For another, suppose that $\mathcal{D}_{P}^{\sharp}$ is ergodic and gapped. and let $\bar{R}$ be the unique density matrix on $\mathcal{H}_{B}$ such that $\mathcal{D}_{P}^{\sharp}\bar{R}=0$. Let $\bar{\rho}_{\gamma}$ be the unique (for $\gamma$ sufficiently large) density matrix on $\mathcal{H}_{AB}$ satisfying $\mathcal{L}_{\gamma}\bar{\rho}_{\gamma}=0$. Then we prove that
@@ 2 Notation and some mathematical preliminaries
Let $\mathcal{H}$ and $\mathcal{K}$ be two finite dimensional Hilbert spaces. Let $\mathcal{T}$ be a linear operator from $\widehat{\mathcal{H}}$ to $\widehat{\mathcal{K}}$; that is a super-operator from $\mathcal{H}$ to $\mathcal{K}$. The adjoint super-operator from $\widehat{\mathcal{K}}$ to $\widehat{\mathcal{H}}$ will always be denoted by $\mathcal{T}^{\dagger}$.
Let $Y$ be an operator on $\mathcal{K}$ and $X$ be an operator on $\mathcal{H}$. Then we may regard $Y$ and $X$ as vectors $\widehat{\mathcal{K}}$ and $\widehat{\mathcal{H}}$ respectively, and define the super-operator $|Y\rangle\langle X|$ from $\widehat{\mathcal{H}}$ to $\widehat{\mathcal{K}}$ by
@@
$\mathcal{P}=\mathcal{P}_{A}\otimes\mathds{1}\ .$ (41)
@@ 4 The super-operators $\mathcal{K}_{P}$ and $\mathcal{D}_{P}$
$\|\mathcal{K}_{P}\|_{1\to 1}\leqslant\|\mathcal{K}\|_{1\to 1}\leqslant 2\|H\|_{\infty}$ (53)
@@
Since $H$ is self-adjoint, (56) can be extended to
$H=\sum_{j=0}^{n_{A}}X_{j}\otimes G_{j}=\sum_{j=0}^{n_{A}}X^{\dagger}_{j}\otimes G^{\dagger}_{j}\ .$ (61)
Then $\mathcal{K}\mathcal{V}R=-i\left(H(\pi_{A}\otimes R)-(\pi_{A}\otimes R)H\right)=-i\sum_{j=0}^{n_{A}}\left(X_{j}\pi_{A}\otimes G_{j}R-\pi_{A}X^{\dagger}_{j}\otimes RG^{\dagger}_{j}\right)$}, and since $\mathcal{S}(Y_{0}\otimes\left(G_{0}R-RG_{0}^{\dagger}\right)=0$, $\mathcal{S}\mathcal{K}(\pi_{A}\otimes R)=-i\sum_{j=1}^{n_{A}}\mathcal{S}\left(X_{j}\pi_{A}\otimes G_{j}R-\pi_{A}X^{\dagger}_{j}\otimes RG^{\dagger}_{j}\right)$}. Using the expression for $\mathcal{S}$ provided in Theorem C.1, namely $\mathcal{S}=\sum_{k=1}^{n_{A}}|Y_{k}\rangle\langle\mathcal{S}^{\dagger}X_{k}|$}.
@@
Using the same expression for $\mathcal{S}$, but this time for the dual pair of bases $\{Y_{0}^{\dagger},\dots,Y_{n_{A}}^{\dagger}\}$ and $\{X_{0}^{\dagger},\dots,X_{n_{A}}^{\dagger}\}$,
@@ 5 Approximate equations of motion
$R(t)=R_{0}+\int_{0}^{t}\mathcal{K}_{P}R(s){\rm d}s+\int_{0}^{t}\textnormal{Tr}_{A}\mathcal{K}\left(\frac{1}{\gamma}\int_{0}^{\gamma s}\left(e^{r\mathcal{D}}\mathcal{Q}\mathcal{K}{\rm d}r\right)Y(s)+\frac{1}{\gamma}\int_{0}^{\gamma s}e^{r\mathcal{D}}\mathcal{Q}\mathcal{K}W(t,r){\rm d}r\right){\rm d}s\ .$ (71)
@@
Again by Theorem C.2 proved in the appendix, for a finite constant $C$ depending only on $\mathcal{D}_{A}$ $\|e^{r\gamma\mathcal{D}}\mathcal{Q}\|_{1\to 1}\leqslant Ce^{-r\gamma a/2}$. Therefore, $\int_{0}^{t}\|F_{1}(s)\|_{1}{\rm d}s\leqslant\frac{4\|\mathcal{K}\|^{2}_{1\to 1}C}{\gamma^{2}a}$}. Combining this with (73) yields $\int_{0}^{t}\|F(s)\|_{1}{\rm d}s\leqslant\frac{C}{\gamma^{2}}$} for a constant $C$ depending only on $\mathcal{D}_{A}$ and $\|\mathcal{K}\|_{1\to 1}$. Now differentiate (74) and apply Lemma 2.1. ∎
@@
The utility of this equation is in doubt. As mentioned in the introduction, an exact calculation in Example 6.1 shows that there are simple models for which $\mathcal{K}_{P}+\gamma^{-1}\mathcal{D}_{P}+\gamma^{-2}\mathcal{B}_{P}$ cannot be put in Lindblad form for any $\gamma>0$. Nonetheless, it still may be that the assumption of an expansion as in (87) could be rigorously justified. Later, we shall prove convergence of a stationary form of the Hilbert expansion for steady state solutions of (1).

### v2

@@ Abstract
We study composite open quantum systems with a finite-dimensional state space $\mathcal{H}_{AB}=\mathcal{H}_{A}\otimes\mathcal{H}_{B}$ governed by a Lindblad equation ${\frac{{\rm d}}{{\rm d}t}\rho(t)=\mathcal{L}_{\gamma}\rho(t)}$ where $\mathcal{L}_{\gamma}\rho=-i[H,\rho]+\gamma\mathcal{D}\rho$. Here, $H$ is a Hamiltonian on $\mathcal{H}_{AB}$ while $\mathcal{D}$ is a dissipator $\mathcal{D}_{A}\otimes\mathds{1}$ acting non-trivially only on part $A$ of the system, which can be thought of as the boundary, and $\gamma$ is a parameter. It is known that the dynamics may simplify as the Zeno limit, $\gamma\to\infty$, is approached, so that after a initial time of order $\gamma^{-1}$, $\rho(t)$ is well approximated by $\pi_{A}\otimes R(t)$ where $\pi_{A}$ is a density matrix on $\mathcal{H}_{A}$ such that $\mathcal{D}_{A}\pi_{A}=0$, and $R(t)$ is an approximate solution of $\frac{{\rm d}}{{\rm d}t}R(t)=\mathcal{L}_{P,\gamma}R(t)$} where $\mathcal{L}_{P,\gamma}R:=-i[H_{P},R]+\gamma^{-1}\mathcal{D}_{P}R$ with $H_{P}$ being a Hamiltonian on $\mathcal{H}_{B}$ and $\mathcal{D}_{P}$ being a Lindblad generator acting on density matrices on $\mathcal{H}_{B}$. We give a rigorous proof of this holding in greater generality than in previous work; we assume only that $\mathcal{D}_{A}$ is ergodic and gapped. Moreover, we precisely control the error terms, and use this to show that the mixing times of $\mathcal{L}_{\gamma}$ and $\mathcal{L}_{P,\gamma}$ are tightly related near the Zeno limit. Despite this connection, the errors in the approximate description of the evolution accumulate on times of order $\gamma^{2}$, so it is difficult to directly access steady states $\bar{\rho}_{\gamma}$ of $\mathcal{L}_{\gamma}$ through study of $\mathcal{L}_{P,\gamma}$. In order to better control the long time behavior, and in particular the steady states $\bar{\rho}_{\gamma}$, we introduce a third Lindblad generator $\mathcal{D}_{P}^{\sharp}$ that does not involve $\gamma$, but is still closely related to $\mathcal{L}_{\gamma}$ and $\mathcal{L}_{P,\gamma}$. We show that if $\mathcal{D}_{P}^{\sharp}$ is ergodic and gapped, then so are $\mathcal{L}_{\gamma}$ and $\mathcal{L}_{P,\gamma}$ for all large $\gamma$, and in this case, if $\bar{\rho}_{\gamma}$ denotes the unique steady state for $\mathcal{L}_{\gamma}$, then $\lim_{\gamma\to\infty}\bar{\rho}_{\gamma}=\pi_{A}\otimes\bar{R}$ where $\bar{R}$ is the unique steady state for $\mathcal{D}_{P}^{\sharp}$. We further show that there is a trace norm convergent expansion $\bar{\rho}_{\gamma}=\pi_{A}\otimes\bar{R}+\gamma^{-1}\sum_{k=0}^{\infty}\gamma^{-k}\bar{n}_{k}$} where, defining $\bar{n}_{-1}:=\pi_{A}\otimes\bar{R}$, $\mathcal{D}\bar{n}_{k}=-i[H,\bar{n}_{k-1}]$ for all $k\geqslant 0$. Using properties of $\mathcal{D}_{P}$ and $\mathcal{D}_{P}^{\sharp}$, we show that this system of equations has a unique solution, and prove convergence. This is illustrated in a simple example for which one can compute $\bar{\rho}_{\gamma}$, and can carry out the expansion explicitly.
@@ 1 Introduction
$\mathcal{D}:=\mathcal{D}_{A}\otimes\mathds{1}_{B}~,$
@@
Without loss of generality, a constant has been added to the Hamiltonian so that $\textnormal{Tr}H=0$ , $H=H_{A}\otimes\mathds{1}_{B}+H_{AB}+\mathds{1}_{A}\otimes H_{B}$ ; here $\mathds{1}_{A}$ and $\mathds{1}_{B}$ are the identity operators on those subspaces, $H_{A}=(1/d_{B})\textnormal{Tr}_{B}H$, $H_{B}=(1/d_{A})\textnormal{Tr}_{A}H$, and $d_{A}$ and $d_{B}$ are the dimensions of $\mathcal{H}_{A}$ and $\mathcal{H}_{B}$.
@@
uniformly in $t\geqslant t_{\gamma}$ , where $R(t)=\textnormal{Tr}_{A}\rho(t)$ is a density matrix on $\mathcal{H}_{B}$. That is, solutions of (1) for a general initial state $\rho_{0}$ and large $\gamma$ rapidly approach $\mathcal{M}$ and then stay close to $\mathcal{M}$, uniformly in time. In Section 3 we give simple proofs of results of this genre in which “closeness” is measured in terms of the trace norm, denoted by $\|\cdot\|_{1}$; see Section 2. In particular, we prove in Theorem 3.1 that for all $\rho_{0}=\pi_{A}\otimes R_{0}\in\mathcal{M}$, there is a constant $C>0$ independent of $\gamma$ such that for all $t>0$,
@@
Note that in (7), the projector $\mathcal{P}$ is applied before $e^{t\mathcal{L}_{\gamma}}$ while in (6), $\mathcal{P}$ is applied after $e^{t\mathcal{L}_{\gamma}}$. These complementary results from Section 3 are important for the study of approximate equations of motion, to which we now turn.
@@
$\lim_{\gamma\to\infty}\|e^{t\mathcal{L}_{\gamma}}\rho_{0}-\pi_{A}\otimes e^{t\mathcal{L}_{P,\gamma}}R_{0}\|_{1}=0\ .$ (17)
Then since for large $\gamma$, $\mathcal{L}_{P,\gamma}$ does not differ much from $\mathcal{K}_{P}$, it follows that (see Theorem 5.3)
@@
In Section 7 we study $e^{-t\mathcal{K}_{P}}e^{t\mathcal{L}_{P,\gamma}}R_{0}$ in the Zeno limit, $\gamma\to\infty$, for general initial data $R_{0}$. We run the initial data forward under the dissipative time evolution governed by $\mathcal{L}_{P,\gamma}$ , and then backwards under the coherent evolution generated by $\mathcal{K}_{P}$ . In the language of scattering theory, we are passing to the “interaction picture” to compare the coherent evolution and its perturbation by a weak dissipation. This is what Davies [13] did in his derivation of Lindblad equations in the weak coupling limit, and hence it is no surprise that his averaging operation introduced there shows up again here.
@@
This has the consequence that if $\mathcal{D}_{P}^{\sharp}$ is ergodic and gapped, then so are $\mathcal{L}_{\gamma}$ and $\mathcal{L}_{P,\gamma}$ for all sufficiently large $\gamma$. See Theorem 7.5, which complements spectral results of [27], for a precise statement.
This shows one way in which $\mathcal{D}_{P}^{\sharp}$ governs the long-time behavior of the evolution described by both $\mathcal{L}_{P,\gamma}$ , and by $\mathcal{L}_{\gamma}$ , near the Zeno limit. For another, suppose that $\mathcal{D}_{P}^{\sharp}$ is ergodic and gapped. and let $\bar{R}$ be the unique density matrix on $\mathcal{H}_{B}$ such that $\mathcal{D}_{P}^{\sharp}\bar{R}=0$. Let $\bar{\rho}_{\gamma}$ be the unique (for $\gamma$ sufficiently large) density matrix on $\mathcal{H}_{AB}$ satisfying $\mathcal{L}_{\gamma}\bar{\rho}_{\gamma}=0$. Then we prove that
@@ 2 Notation and some mathematical preliminaries
Let $\mathcal{H}$ and $\mathcal{K}$ be two finite dimensional Hilbert spaces. Let $\mathcal{T}$ be a linear operator from $\widehat{\mathcal{H}}$ to $\widehat{\mathcal{K}}$; that is, a super-operator from $\mathcal{H}$ to $\mathcal{K}$. The adjoint operator from $\widehat{\mathcal{K}}$ to $\widehat{\mathcal{H}}$ (a super operator from $\mathcal{K}$ to $\mathcal{H}$) will always be denoted by $\mathcal{T}^{\dagger}$.
Let $Y$ be an operator on $\mathcal{K}$ and $X$ be an operator on $\mathcal{H}$. Then we may regard $Y$ and $X$ as vectors in the Hilbert spaces $\widehat{\mathcal{K}}$ and $\widehat{\mathcal{H}}$ respectively, and define the super-operator $|Y\rangle\langle X|$ from $\widehat{\mathcal{H}}$ to $\widehat{\mathcal{K}}$ by
@@
$\mathcal{P}=\mathcal{P}_{A}\otimes\mathds{1}\ .$
@@ 4 The super-operators $\mathcal{K}_{P}$ and $\mathcal{D}_{P}$
$\|\mathcal{K}_{P}\|_{1\to 1}\leqslant\|\mathcal{K}\|_{1\to 1}\leqslant 2\|H\|_{\infty}$
@@
Since $H$ is self-adjoint, (53) can be extended to $H=\sum_{j=0}^{n_{A}}X_{j}\otimes G_{j}=\sum_{j=0}^{n_{A}}X^{\dagger}_{j}\otimes G^{\dagger}_{j}$}. Using both of these expression for $H$,
$\mathcal{K}\mathcal{V}R=-i\left(H(\pi_{A}\otimes R)-(\pi_{A}\otimes R)H\right)=-i\sum_{j=0}^{n_{A}}\left(X_{j}\pi_{A}\otimes G_{j}R-\pi_{A}X^{\dagger}_{j}\otimes RG^{\dagger}_{j}\right)\ ,$
and since $\mathcal{S}(Y_{0}\otimes\left(G_{0}R-RG_{0}^{\dagger}\right)=0$, $\mathcal{S}\mathcal{K}(\pi_{A}\otimes R)=-i\sum_{j=1}^{n_{A}}\mathcal{S}\left(X_{j}\pi_{A}\otimes G_{j}R-\pi_{A}X^{\dagger}_{j}\otimes RG^{\dagger}_{j}\right)$}. Using the expression for $\mathcal{S}$ provided in Theorem C.1, namely $\mathcal{S}=\sum_{k=1}^{n_{A}}|Y_{k}\rangle\langle\mathcal{S}^{\dagger}X_{k}|$}.
@@
Using the same expression for $\mathcal{S}$, but now for the dual bases $\{Y_{0}^{\dagger},\dots,Y_{n_{A}}^{\dagger}\}$ and $\{X_{0}^{\dagger},\dots,X_{n_{A}}^{\dagger}\}$,
@@ 5 Approximate equations of motion
$R(t)=R_{0}+\int_{0}^{t}\mathcal{K}_{P}R(s){\rm d}s+\int_{0}^{t}\textnormal{Tr}_{A}\mathcal{K}\left(\frac{1}{\gamma}\int_{0}^{\gamma s}\left(e^{r\mathcal{D}}\mathcal{Q}\mathcal{K}{\rm d}r\right)Y(s)+\frac{1}{\gamma}\int_{0}^{\gamma s}e^{r\mathcal{D}}\mathcal{Q}\mathcal{K}W(t,r){\rm d}r\right){\rm d}s\ .$
@@
Again by Theorem C.2 proved in the appendix, for a finite constant $C$ depending only on $\mathcal{D}_{A}$ $\|e^{r\gamma\mathcal{D}}\mathcal{Q}\|_{1\to 1}\leqslant Ce^{-r\gamma a/2}$. Therefore, $\int_{0}^{t}\|F_{1}(s)\|_{1}{\rm d}s\leqslant\frac{4\|\mathcal{K}\|^{2}_{1\to 1}C}{\gamma^{2}a}$}. Combining this with (68) yields $\int_{0}^{t}\|F(s)\|_{1}{\rm d}s\leqslant\frac{C}{\gamma^{2}}$} for a constant $C$ depending only on $\mathcal{D}_{A}$ and $\|\mathcal{K}\|_{1\to 1}$. Now differentiate (69) and apply Lemma 2.1 and (34). ∎
@@
The utility of this equation is in doubt. As mentioned in the introduction, an exact calculation for the model in Example 6.1 shows that there are simple cases in which $\mathcal{K}_{P}+\gamma^{-1}\mathcal{D}_{P}+\gamma^{-2}\mathcal{B}_{P}$ cannot be put in Lindblad form for any $\gamma>0$. Nonetheless, it still may be that the assumption of an expansion as in (82) could be rigorously justified. Later, we shall prove convergence of a stationary form of the Hilbert expansion for steady state solutions of (1).